# MCCoder: Streamlining Motion Control with LLM-Assisted Code Generation and Rigorous Verification

### v2

@@ II-B Control Code and Motion Library
III METHODOLOGY
@@ III-B Hybrid Retrieval
III-C Control Code Generation
@@ IV-A Construction
To replicate a human engineer’s problem-solving skills, we categorized the tasks into three difficulty levels, with examples provided in Section IV-A. Difficulty Level 1 involves calling standard APIs with straightforward arguments as outlined in the documentation. Difficulty Level 2 features APIs with more complex parameter types, including easily confused overrides, as well as potential user errors such as typos or improper parameter instructions. LLMs may err here but should leverage error feedback or retrieved content for self-correction. Difficulty Level 3 requires integrating multiple motion tasks, not merely stacking simple ones, but combining them into a cohesive set linked by events and dependencies. Solving these demands robust retrieval of all subtasks and a comprehensive understanding of the overall task. These levels mirror real-world motion control scenarios, progressing from simple to complex.
@@ IV-B Metrics
IV-B1 First Time Pass Rate (FTPR)
@@ IV-B2 MatchEndPoints and DTW
V EXPERIMENTS
V-A Experimental Settings
@@ V-D Experimental Results
V-D1 RQ1. Does MCCoder improve the code generation performance?
@@
V-D2 RQ2. Which base model behaves the best with MCCoder?
@@
V-D3 RQ3. Why is MCCoder’s data verification critical for motion control?
@@ V-E Error Analysis
VI DISCUSSION
VI-A Limitations and Future Work
@@
Furthermore, deploying MCCoder in real-world production environments necessitates fine-tuned, edge-side small models tailored to proprietary motion libraries. Integrating these specialized models with RAG techniques will significantly enhance data security and mitigate model hallucinations, making this an important research direction moving forward.
VII RELATED WORK
VII-A Code Generation with LLMs and Strategies
VII-B Code Generation in Motion Control

### v3

@@ II-B Control Code and Motion Library
III RELATED WORK
III-A Code Generation with LLMs and Strategies
III-B Code Generation in Motion Control
IV METHODOLOGY
@@ IV-B Hybrid Retrieval
IV-C Control Code Generation
@@ V-A Construction
To replicate a human engineer’s problem-solving skills, we categorized the tasks into three difficulty levels, with examples provided in Section V-A. Difficulty Level 1 involves calling standard APIs with straightforward arguments as outlined in the documentation. Difficulty Level 2 features APIs with more complex parameter types, including easily confused overrides, as well as potential user errors such as typos or improper parameter instructions. LLMs may err here but should leverage error feedback or retrieved content for self-correction. Difficulty Level 3 requires integrating multiple motion tasks, not merely stacking simple ones, but combining them into a cohesive set linked by events and dependencies. Solving these demands robust retrieval of all subtasks and a comprehensive understanding of the overall task. These levels mirror real-world motion control scenarios, progressing from simple to complex.
@@ V-B Metrics
V-B1 First Time Pass Rate (FTPR)
@@ V-B2 MatchEndPoints and DTW
VI EXPERIMENTS
VI-A Experimental Settings
@@ VI-D Experimental Results
VI-D1 RQ1. Does MCCoder improve the code generation performance?
@@
VI-D2 RQ2. Which base model behaves the best with MCCoder?
@@
VI-D3 RQ3. Why is MCCoder’s data verification critical for motion control?
@@ VI-E Error Analysis
Self-correction is designed to address these returned errors efficiently. For general syntax errors, the LLM attempts direct self-repair. When errors involve APIs or arguments and originate from the Soft-Motion engine, the LLM further leverages error messages as references by searching related documentation to guide correction. These processes are typically fast, but to ensure time efficiency and smooth operation of the entire MCCoder system, the number of self-correction attempts is deliberately limited, for example, to three retries.
VII DISCUSSION
VII-A Limitations and Future Work
@@
A limitation of the current work is the lack of evaluation on more models, especially smaller-parameter ones, and the absence of systematic assessment of whether a well-designed workflow could enable small models to approach large-model performance. However, deploying MCCoder in real-world industrial field environments necessitates fine-tuned, edge-side small models tailored to proprietary motion libraries. These small models generate code faster than cloud-based large models and offer stronger data security, which is crucial on site. Fine-tuning ensures solid basic automation programming capabilities, while RAG allows dynamic incorporation of the latest code, templates, or user-defined patterns, achieving large-model-level programming advancements and further reducing hallucinations—making this a key research direction.